OpenStudy (anonymous):

A geometric series has a sum of 9,837. The common ratio is 3 and the first term is 9. How many terms are in the series?

OpenStudy (dumbcow):

\[s _{n} = \frac{a(r^{n}-1)}{r-1}\] s = 9837 a = 9 r = 3 solve for n

OpenStudy (anonymous):

use the formula Sn = a * (r^n - 1)/ (r - 1) n = no. terms r = common ratio and a = ist term

OpenStudy (anonymous):

9?

OpenStudy (anonymous):

7

OpenStudy (anonymous):

dedfinitely 7 - sorry math must go